# Emergence of Newtonian Deterministic Causality from Stochastic Motions in Continuous Space and Time

### v1

@@ Abstract
Since Newton’s time, deterministic causality has been considered a crucial prerequisite in any fundamental theory in physics. In contrast, the present work investigates stochastic dynamical models for motion in one spatial dimension, in which Newtonian mechanics becomes an emergent property: We present a coherent theory in which a Hamilton-Jacobi equation (HJE) emerges in a description of the evolution of entropy $-\phi(x,t)=\epsilon\log$(Probability) of a system under observation and in the limit of large information extent $\epsilon^{-1}$ in homogeneous space and time. The variable $\phi$ represents a non-random high-order statistical concept that is distinct from probability itself as $\epsilon=0$; the HJE embodies an emergent law of deterministic causality in continuous space and time with an Imaginary Scale symmetry $(t,x,\phi)\leftrightarrow(it,ix,i\phi)$. $\phi(x,t)$ exhibits a nonlinear wave phenomenon with a mathematical singularity in finite time, overcoming which we introduce viscosity $\epsilon(\partial^{2}\phi/\partial x^{2})$ and wave $i\epsilon(\partial^{2}\phi/\partial x^{2})$ perturbations, articulating dissipation and conservation, which break the Imaginary Scale symmetry: They lead to the Brownian motion and Schrödinger’s equation of motion, respectively. Last but not least, Lagrange’s action in classical mechanics acquires an entropic interpretation and Hamilton’s principle is established.
@@ I Introduction
The profound insights from statistical physics, particularly in the theory of phase transition, have provided us with an intriguing recipe for generating new principles in the age of machine learning via the idealization of data ad infinitum Yang and Qian (2022). In the traditional statistical thermodynamics, the “infinity” has always been identified with the largeness of a macroscopic system, whose equilibrium possesses an Eulerian homogeneous degree-$1$ entropy function $S(E,V,N)$ Callen (1991). In a recent series of papers on the thermodynamics of small systems Qian (2022); Lu and Qian (2022); Qian (2021), it was suggested that T. L. Hill’s theory Hill (1964); Bedeaux et al. (2023) should be framed as for repeated measurements in time, with “infinity” being identified with a scientific measurement over a $\Delta t$ which, no matter how small, nevertheless is very large in a continuous time that is infinitely divisible.
@@
We consider the limit of random motions in continuous space and time in the present work. The authors are aware of the fact that in mathematical literature, the notion of random processes in continuous space and time has been rigorously defined in a much more refined way via Brownian motions Wang and Uhlenbeck (1945) and Lévy processes Khoshnevisan and Schilling (2016). From the perspective of a complex world at $10^{-100}$ meter and $10^{-100}$ second whose existence we postulate, any scientific $\Delta x$ and $\Delta t$ are extensive quantities whose ratio, $\Delta x/\Delta t$, defines a “macroscopic” velocity. We seek emergent thermodynamic law(s) in this mathematical idealization and find a Hamilton-Jacobi equation (HJE) for entropy, shown in Eq. (3), which arises in this context. In classical physics, HJE is an integral part of Newtonian mechanics Landau and Lifshitz (1976); Evans (2010) that obeys strictly the law of causality with determinism. The success of Newtonian mechanics has made many people to believe in their subconsciousness that the Principle of Deterministic Causality (PDC) should be a built-in fundamental in any physics theory Born (1949); Monod (1971); Ruelle (1993); Hehl (2021). The development of the special theory of relativity seemed to have strengthened this belief de Broglie (1964).

### v2

@@ Abstract
Since Newton’s time, deterministic causality has been considered a crucial prerequisite in any fundamental theory in physics. In contrast, the present work investigates stochastic dynamical models for motion in one spatial dimension, in which Newtonian mechanics becomes an emergent property: We present a coherent theory in which a Hamilton-Jacobi equation (HJE) emerges in a description of the evolution of entropy $-\phi(x,t)=\epsilon\log$(Probability) of a system under observation and in the limit of large information extent $\epsilon^{-1}$ in homogeneous space and time. The variable $\phi$ represents a non-random high-order statistical concept that is distinct from probability itself as $\epsilon=0$; the HJE embodies an emergent law of deterministic causality in continuous space and time with an Imaginary Scale symmetry $(t,x,\phi)\leftrightarrow(it,ix,-i\phi)$. $\phi(x,t)$ exhibits a nonlinear wave phenomenon with a mathematical singularity in finite time, overcoming which we introduce viscosity $\epsilon(\partial^{2}\phi/\partial x^{2})$ and wave $i\epsilon(\partial^{2}\phi/\partial x^{2})$ perturbations, articulating dissipation and conservation, which break the Imaginary Scale symmetry: They lead to the Brownian motion and Schrödinger’s equation of motion, respectively. Last but not least, Lagrange’s action in classical mechanics acquires an entropic interpretation and Hamilton’s principle is established.
@@ I Introduction
The profound insights from statistical physics, particularly in the theory of phase transition, have provided us with an intriguing recipe for generating new principles in the age of machine learning via the idealization of data ad infinitum Yang and Qian (2022, 2024). In the traditional statistical thermodynamics, the “infinity” has always been identified with the largeness of a macroscopic system, whose equilibrium possesses an Eulerian homogeneous degree-$1$ entropy function $S(E,V,N)$ Callen (1991). In a recent series of papers on the thermodynamics of small systems Qian (2022); Lu and Qian (2022); Qian (2024, 2021), it was suggested that T. L. Hill’s theory Hill (1964); Bedeaux et al. (2023) should be framed as for repeated measurements in time, with “infinity” being identified with a scientific measurement over a $\Delta t$ which, no matter how small, nevertheless is very large in a continuous time that is infinitely divisible.
@@
We consider the limit of random motions in continuous space and time in the present work. The authors are aware of the fact that in mathematical literature, the notion of random processes in continuous space and time has been rigorously defined in a much more refined way via Brownian motions Wang and Uhlenbeck (1945) and Lévy processes with their Lev́y-Khinchin characteristic exponents Khoshnevisan and Schilling (2016). From the perspective of a complex world at $10^{-100}$ meter and $10^{-100}$ second whose existence we postulate, any scientific $\Delta x$ and $\Delta t$ are extensive quantities whose ratio, $\Delta x/\Delta t$, defines a “macroscopic” velocity. We seek emergent thermodynamic law(s) in this mathematical idealization and find a Hamilton-Jacobi equation (HJE) for entropy, shown in Eq. (3), which arises in this context. In classical physics, HJE is an integral part of Newtonian mechanics Landau and Lifshitz (1976); Evans (2010) that obeys strictly the law of causality with determinism. The success of Newtonian mechanics has made many people to believe in their subconsciousness that the Principle of Deterministic Causality (PDC) should be a built-in fundamental in any physics theory Born (1949); Monod (1971); Ruelle (1993); Hehl (2021). The development of the special theory of relativity seemed to have strengthened this belief de Broglie (1964).
@@ Appendix C Derivation of Entropy Evolution Equation (3)
Appendix D The physics of continuous space and time limit
As $\Delta x$, $\Delta t\to 0$, Newtonian smooth motion simply assumes the existence of $\Delta x/\Delta t=v$. Stochastic motion has increment in $\Delta t$ time, $X(t+\Delta t)-X(t)=\Delta x$, $-\Delta x$, 00 with probabilities $p$, $q$, $r$, and $p+q+r=1$. The canonical diffusion limit supposes the existence of
$\displaystyle\frac{(p-q)\Delta x}{\Delta t}=V,\ \frac{(1-r)(\Delta x)^{2}}{2\Delta t}=D,\text{ and}$
$\displaystyle\frac{1}{\Delta x}\mathbb{P}\big{\{}X(t)\in[\,x,x+\Delta x\,]\,\big{\}}=f(x,t),$
as $\Delta t\to 0$. This imply $\frac{2D(p-q)^{2}}{(1-r)V^{2}}\sim\Delta t$. If, however, $p,q,r$ are independent of $\Delta t$, then the existence of $D$ implies $\frac{\Delta x}{\Delta t}\to\infty$. The resulting overdamped Brownian motion $B(t)$ is nowhere differentiable Durrett (2019).
In contradistinction the present work supposes the existence of
$\displaystyle\frac{\Delta x}{\epsilon}=\eta,\ \frac{\Delta t}{\epsilon}=\gamma^{-1},\text{ and}$
$\displaystyle-\epsilon\log\mathbb{P}\big{\{}X_{\epsilon}(t)\in{\rm d}x\big{\}}=\phi(x,t),$
as $\epsilon\to 0$; thus $\Delta x/\Delta t=\gamma\eta$ exists. We are interested in the motion on $\epsilon^{-1}$ scale; the $\phi$ satisfies an HJE with the Hamiltonian function given in Eq. (61). The corresponding Hamiltonian dynamics then is
$\dot{x}=\frac{\gamma\eta\big{(}pe^{\eta y}-qe^{-\eta y}\big{)}}{pe^{\eta y}+r+qe^{-\eta y}},\ \ \dot{y}=0.$ (64)
$\dot{x}$ given above is a monotonic function of $y$ and $|\dot{x}|\to\gamma\eta$ as $|y|\to\infty$. One now has $|\dot{x}|\leq\gamma\eta$, a limit that is independent from $p,q,r$. This is the meaning of undamped/conserved Hamiltonian motion.
Introducing $\tilde{v}\equiv\dot{x}/(\gamma\eta)$, and solving $y$ in terms of $\tilde{v}$ from (64), $|\tilde{v}|\leq 1$:
$y=\frac{1}{\eta}\log\left\{\frac{r\tilde{v}+\sqrt{4pq-(4pq-r^{2})\tilde{v}^{2}}}{2p\big{(}1-\tilde{v}\big{)}}\right\}.$ (65)
For isotropic space with $p=q=(1-r)/2$, $y(\tilde{v})$ is an odd function. As $\tilde{v}\to 0$, $y\simeq\tilde{v}/(\eta(1-r))$ and as $|\tilde{v}|\to 1$,
$y$ $=$ $\frac{1}{2\eta}\log\left\{\frac{1-\tilde{v}}{1+\tilde{v}}\,\frac{\sqrt{r^{2}+(1-2r)(1-\tilde{v}^{2})}-r\tilde{v}}{\sqrt{r^{2}+(1-2r)(1-\tilde{v}^{2})}+r\tilde{v}}\right\}$
$\simeq$ $\frac{1}{2\eta}\log\left\{\left(\frac{1-\tilde{v}}{1+\tilde{v}}\right)^{2}\frac{2r^{2}+(1-2r)(1+\tilde{v})}{2r^{2}+(1-2r)(1-\tilde{v})}\right\}$
$\simeq$ $\frac{1}{\eta}\log\left(\frac{1-\tilde{v}}{1+\tilde{v}}\right)$
$+$ $\frac{\tilde{v}}{\eta|\tilde{v}|}\left\{\log\frac{1-r}{r}-a_{1}\big{(}1-|\tilde{v}|\big{)}+a_{2}\big{(}1-|\tilde{v}|\big{)}^{2}\right\},$
where $a_{1}=\frac{(1-2r)(1-2r+2r^{2})}{4r^{2}(1-r)^{2}}$ and $a_{2}=\frac{(1-2r)^{2}[(1-r)^{4}-r^{4}]}{16r^{4}(1-r)^{4}}$.